a mans age is three times that of his son.twelve year's ago the man's age was 27 times his sons age.How old are they now?
Answer by lwsshak3(11628)

a mans age is three times that of his son.twelve year's ago the man's age was 27 times his sons age.How old are they now?
***
let x=son's age
3x=father's age
..
12 yrs ago:
x-12=son's age
3x-12=father's age
father's age=27 times son's age
3x-12=27(x-12)
3x-12=27x-324
24x=312
x=13
3x=39
son's age=13
father's age=39